# Online Multi-Task Offloading for Semantic-Aware Edge Computing Systems

### v1

@@ III Multi-Agent Proximal Policy Optimization for Resource Allocation
The binary offloading variable in Constraint (12b) introduces non-convexity, and the QoE is characterized by a non-convex logistic function, rendering the problem P0 a non-convex MINLP problem. To solve the non-convex MINLP problem P0, we employ an advanced MAPPO algorithm, an actor-critic-based reinforcement learning algorithm, which is suitable for both discrete and continuous action spaces. To utilize the MAPPO algorithm, we reformulate problem P0 as an MDP problem and detail the training process of the MAPPO.
@@ III-A Problem Reformulation Based on MDP
A 4-tuples MDP $\langle\mathcal{S},\mathcal{A},\mathcal{R},\gamma\rangle$ is used to describe the interaction between the ES and the wireless network. where $\mathcal{S}$, $\mathcal{A}$, $\mathcal{R}$ and $\gamma\in[0,1)$ denote the state space, action space, reward function, and the discount factor, respectively. As agents, UEs need to collect observable states to input into the policy network and obtain their actions. After executing actions, UEs update their observable states and the policy network based on corresponding rewards to improve the output of actions. Specifically, the observable state of UEn can be represented as $o_{n}=\{g_{k,n},l_{n}^{U},l_{n}^{S}\}\in\mathcal{S}$. The action space of UEn refers to the variables that need to be optimized, which can be expressed as $a_{n}=\{\rho_{n},p_{n},f_{n},\mu_{n},x_{k,n}\}\in\mathcal{A}$. UEs need to continuously optimize offloading strategies, transmit power, local computation frequency, semantic compression ratio, and channel assignment to purchase better QoE. Drawing from the optimization problem P0, it is evident that the reward function needs to focus on the summation QoE of all UEs. In addition to QoE, the reward function includes penalty terms for latency and task accuracy, expressed as
@@
Therefore, the total reward at the $t$th training timestep is defined as $r_{t}=\sum^{N}_{n=1}r_{n}$. The cumulative long-term reward to account for the impact of the current state and action on the expected future rewards:
@@
$J\left(\boldsymbol{\pi}_{\boldsymbol{\theta}}\right)=\mathbb{E}_{\boldsymbol{a}_{t}\sim\pi_{\boldsymbol{\theta}},\boldsymbol{s}_{t}\sim\mathcal{P}}\left[\sum_{t=1}^{T}\gamma^{t-1}r_{t}\left(\boldsymbol{s}_{t},\boldsymbol{a}_{t}\right)\right].$ (15)
@@
III-B PPO-Based Algorithm
For on-policy gradient algorithms, trajectories are typically sampled using a fixed policy $\pi_{\boldsymbol{\theta}}$, and policy updates are made using only those trajectories sampled by $\pi_{\boldsymbol{\theta}}$. Fig. 4 illustrates our proposed PPO-based algorithm. PPO is an off-policy algorithm that improves sampling efficiency significantly. It utilizes an experience replay buffer, which stores trajectories from past time steps. Policy $\pi_{\boldsymbol{\theta}}$ randomly selects trajectories from the replay buffer for updates, enhancing sampling efficiency. This implementation benefits from the use of importance sampling, as outlined in Lemma 1.
@@ III-B PPO-Based Algorithm
$Q_{t}^{\boldsymbol{\theta}^{\prime}}\left(\boldsymbol{s}_{t},\boldsymbol{a}_{t}\right)=\mathbb{E}_{\boldsymbol{a}_{t}\sim\pi_{\boldsymbol{\theta}},\boldsymbol{s}_{t}\sim\mathcal{P}}\left[\sum_{t=1}^{T}\gamma^{t-1}r_{t}\right],$ (20)
@@
$V_{t}^{\boldsymbol{\theta}^{\prime}}\left(\boldsymbol{s}_{t}\right)=\mathbb{E}_{\boldsymbol{s}_{t}\sim\mathcal{P}}\left[\sum_{t=1}^{T}\gamma^{t-1}r_{t}\right].$ (21)
@@
where $\delta_{t}=A_{t}^{\boldsymbol{\theta}^{\prime}}\left(\boldsymbol{s}_{t},\boldsymbol{a}_{t}\right)$. Note that (23) aims to balance the variance and bias of the advantage estimates by leveraging information from multiple future steps while discounting their influence.
Figure 4: The framework of PPO-based algorithm.
If there are no constraints on the policies $\pi_{\boldsymbol{\theta}}$ and $\pi_{\boldsymbol{\theta}^{\prime}}$, significant fluctuations may occur during policy updates. This is due to the limited sampling which fails to accurately represent the probability distributions described in Lemma 1. To constrain the divergence between policies $\pi_{\boldsymbol{\theta}}$ and $\pi_{\boldsymbol{\theta}^{\prime}}$, $J^{CLIP}(\boldsymbol{\theta})$ is employed. It is defined as
@@
$=\mathbb{E}_{t}\left[\min(r_{t}(\boldsymbol{\theta})A_{t}^{\boldsymbol{\theta}^{\prime}}\left(\boldsymbol{s}_{t},\boldsymbol{a}_{t}\right),g(\epsilon,A_{t}^{\boldsymbol{\theta}^{\prime}}\left(\boldsymbol{s}_{t},\boldsymbol{a}_{t}\right)))\right],$
@@ IV Simulation Results
In this section, the performance of the semantic-aware multi-modal task offloading model will be demonstrated. We provide specific implementation details, investigate the QoE performance of the proposed approach, and examine the impact of user preferences on the system.
@@ IV-A Implementation Details
Three tasks are considered: text classification, image classification, and VQA. For text classification, we use the SST-2 dataset, for image classification, the CIFAR-10 dataset, and for VQA, the VQAv2 dataset is adopted. The text semantic extraction network utilizes an advanced BERT-based [45] text embedding model, while the image semantic extraction network is based on the Vision Transformer (ViT) [46]. Training is conducted using the AdamW optimizer, with a learning rate of $3\times 10^{-5}$, a batch size of 50, and a weight decay of $1\times 10^{-4}$. All experiments were conducted on a platform equipped with GPU NVIDIA Titan V and CPU Intel i9-9900K @3.6 GHz. When the training of the multi-modal semantic transmission model is complete, we can calculate the GPU consumption for basic semantic extraction $l_{n}^{SE}$ and the GPU consumption $l^{S}$ for execution at the ES. Typically, the computational demand of the model deployed at the ES is greater than that of the local model, which leads to better task performance. Therefore, we set $l_{n}^{U}$ (local execution GPU consumption) to be less than $l^{S}$ (ES execution GPU consumption), with the task accuracy of local execution being lower than that of execution at the ES.
We consider scenarios with two users and two orthogonal subcarriers, where each user has a queue length of 10, containing tasks from different modalities. Additional details regarding the wireless environment, GPU computational frequency, and MAPPO training configurations are summarized in Table II. To verify the effectiveness of the semantic-aware multi-modal task offloading framework, we consider the following benchmarks:
@@
Number of UEs $N$ 2
Carrier frequency 2 GHz
Sub-channel bandwidth $B$ 2 MHz
Channel gain $g_{k,n}$ range (0.1, 1)
@@
Transmit power $p_{n}$ range (10, 100) mW
Semantic extraction factor $\mu_{n}$ (0.3, 1)
FLOPS per cycle of local GPU $o_{n}$ 2
Local GPU frequency $f_{n}$ range (0.55, 1.0) GHz
FLOPS per cycle of remote GPU $o_{s}$ 4
Remote GPU frequency $f_{s}$ 3 GHz
@@
Energy consumption of ES $\alpha^{S}$ $5\times 10^{-27}$
Minimum accuracy requirements $\epsilon_{\min}$ 0.56
@@
Batch size of training the MAPPO 1
@@
Learning rate $5\times 10^{-4}$
Advantage discount factor $\lambda$ 0.98
Reward discount factor $\gamma$ 0.95
@@
Policy entropy bonus weight $b_{2}$ 0.01
  • Exhaustive search: This method explores the entire action space to find the upper bound of the optimization problem P0. The action space is discretized to facilitate finding the optimal discrete solution.
  • DQN: DQN can only handle discrete action spaces, so we discretized the action space and used the same environment as MAPPO.
  • Semantic-unaware: This benchmark transmits all data without semantic extraction during offloading, i.e. $\mu_{n}=1$. All other settings are identical to those used in MAPPO.
@@ IV-B QoE Performance and Convergence
1) Convergence Performance: The preference parameters, $\omega_{t}$, $\omega_{e}$, and $\omega_{a}$ are all set to 1/3, indicating that the model has no specific preference among latency, energy consumption, and accuracy. Fig. 6 shows that the MAPPO algorithm converges after approximately 50 episodes and maintains optimal performance. For the DQN algorithm, the replay buffer is set to 50 episodes, which causes a jump at 50 episodes in Fig. 7. The model finally converges after 350 episodes.
TABLE III: QoE Performance without Preference.
Local DQN MAPPO Semantic-Unaware MAPPO Exhaustive Search
QoE 0.5 0.546 0.6 0.581 0.604
2) QoE Performance: Table. III presents the QoE performance without preference. All methods were given the same task queue and channel conditions as the exhaustive search to ensure fairness in comparisons. Table. III shows that the MAPPO method closely approaches the exhaustive search, demonstrating that our model performs near the optimal level. The semantic-unaware MAPPO method fixes $\mu_{n}=1$ to transmit all data without any compression, resulting in increased transmission latency and reduced QoE compared to the MAPPO method. Conventional reinforcement learning approaches like DQN perform only slightly better than local execution. This is because DQN operates within a discrete action space and solely optimizes the value function. The proposed MAPPO method optimizes both the policy function and the value function, supporting a continuous action space, and thereby achieving near-optimal QoE. Local execution is solely related to the local computation frequency. When calculating QoE, the local computation frequency is fixed, meaning we use the local execution’s latency, energy consumption, and accuracy to normalize the latency, energy, and accuracy obtained from other methods. Thus, the QoE of local execution is fixed at 0.5.
@@
Figure 7: Convergence of the DQN Reward.
To further investigate the contributions of latency, energy consumption, and task performance to QoE, Fig. 8 illustrates the execution latency when the task queue consists of mixed tasks. To eliminate the influence of task type distribution on execution latency, we tested 100 rounds and averaged the execution latency. Fig 8 shows that the MAPPO method has a lower task execution latency compared to the benchmarks. Although the DQN algorithm’s overall QoE is lower than that of the semantic-unaware MAPPO method, its performance in execution latency is superior. This indicates that DQN’s optimization focuses more on reducing execution latency. Then we fix the task queue for all UEs to consist of either text classification, image classification, or VQA, and test across 100 rounds to calculate the average. Fig. 9 further reveals the execution latency for different tasks. Text classification involves less computation than image classification and VQA, resulting in lower execution latency. It is evident that MAPPO achieves the lowest execution latency for all tasks. Local execution is the slowest across all task types. DQN and the semantic-unaware MAPPO method show very similar execution latencies for text classification tasks, while each has its strengths and weaknesses in image classification and VQA tasks.
In addition to execution latency, QoE also includes energy consumption and task performance. Table IV provides the energy consumption and task accuracy for all different schemes. Local execution has the lowest energy consumption, but its task accuracy is also the lowest. This is because local GPUs have limited computing power, and despite the lower energy consumption advantage compared to ES, it doesn’t excel in task performance. The semantic-unaware method achieves the best task performance because it transmits significantly more data than the MAPPO method. However, this also results in higher energy consumption compared to the MAPPO method.
TABLE IV: Energy Consumption and Task Performance without Preference
Local DQN MAPPO Semantic-Unaware MAPPO
Energy Consumption (J) 1.37 3.05 3.19 3.60
Accuracy (%) 58.27 61.05 64.37 69.15
Figure 8: Delay performance of mix tasks.
(a) Execution time on text classification.
(b) Execution time on image classification.
(c) Execution time on VQA.
Figure 9: Delay performance of different task modalities. Figs. (a), (b), and (c) depict the user task queue set for text classification, image classification, or VQA, respectively, and measure the task execution time for different methods.

### v2

@@ III Multi-Agent Proximal Policy Optimization for Resource Allocation
The binary offloading variable in Constraint (11b) introduces non-convexity, and the QoE is characterized by a non-convex logistic function, rendering the problem P0 a non-convex MINLP problem. To solve the non-convex MINLP problem P0, we propose an advanced semantic-aware MAPPO algorithm, i.e., an actor-critic-based reinforcement learning algorithm, which is suitable for both discrete and continuous action spaces. To utilize the semantic-aware MAPPO algorithm, we reformulate problem P0 as an MDP and detail the training process of the MAPPO.
@@ III-A Problem Reformulation Based on MDP
A 4-tuple MDP $\langle\mathcal{S},\mathcal{A},\mathcal{R},\gamma\rangle$ is used to describe the interaction between the ES and the wireless network, where $\mathcal{S}$, $\mathcal{A}$, $\mathcal{R}$ and $\gamma\in[0,1)$ denote the state space, action space, reward function, and the discount factor, respectively. As agents, UEs need to collect observable states to input into the policy network and obtain their actions. After executing actions, UEs update their observable states and the policy network based on corresponding rewards to improve the output of actions. Specifically, the observable state of UEn can be represented as $o_{n}=\{g_{k,n},l_{n}^{U},l_{n}^{S}\}\in\mathcal{S}$. The action space of UEn refers to the variables that need to be optimized, which can be expressed as $a_{n}=\{\rho_{n},p_{n},f_{n},\mu_{n},x_{k,n}\}\in\mathcal{A}$. UEs need to continuously optimize offloading strategies, transmit power, local computation frequency, semantic compression ratio, and channel assignment to purchase better QoE. Drawing from the optimization problem P0, it is evident that the reward function needs to focus on the summation QoE of all UEs. In addition to QoE, the reward function includes penalty terms for latency and task accuracy, expressed as
@@
Therefore, the total reward at the $t$th training timestep is defined as $r_{t}=\sum^{N}_{n=1}r_{n}$. The cumulative long-term reward accounts for the impact of the current state and action on the expected future rewards:
@@
$J\left(\pi_{\boldsymbol{\theta}}\right)=\mathbb{E}_{\boldsymbol{a}_{t}\sim\pi_{\boldsymbol{\theta}},\boldsymbol{s}_{t}\sim\mathcal{P}}\left[\sum_{t=1}^{T}\gamma^{t-1}r_{t}\left(\boldsymbol{s}_{t},\boldsymbol{a}_{t}\right)\right].$ (14)
@@
III-B Semantic-aware MAPPO Algorithm
Fig. 4 illustrates our proposed MAPPO algorithm. PPO is an on-policy algorithm whose trajectories are typically sampled using a fixed policy $\pi_{\boldsymbol{\theta}}$, and policy updates are made using only those trajectories sampled by $\pi_{\boldsymbol{\theta}}$. It utilizes an experience replay buffer, which stores trajectories from past time steps. To enhance environmental exploration efficiency, the MAPPO algorithm employs a parallelized architecture with multiple threads that simultaneously collect environment interaction data and synchronize it during updates. To enhance data efficiency, the MAPPO algorithm typically uses small batches of data for multi-step updates. This leads to a slight discrepancy between the current policy and the policy used during data generation during the update process. The use of importance sampling techniques allows for the assessment of this policy deviation and adjusts the gradient updates, thus stabilizing the learning process. The importance sampling is outlined in Lemma 1.
@@ III-B Semantic-aware MAPPO Algorithm
$Q_{t}^{\boldsymbol{\theta}^{\prime}}\left(\boldsymbol{s}_{t},\boldsymbol{a}_{t}\right)=\mathbb{E}_{\boldsymbol{a}_{t}\sim\pi_{\boldsymbol{\theta^{\prime}}},\boldsymbol{s}_{t}\sim\mathcal{P}}\left[\sum_{t^{\prime}=t}^{T}\gamma^{t^{\prime}-t}r_{t^{\prime}}\right],$ (19)
@@
$V_{t}^{\boldsymbol{\theta}^{\prime}}\left(\boldsymbol{s}_{t}\right)=\mathbb{E}_{\boldsymbol{s}_{t}\sim\mathcal{P}}\left[\sum_{t^{\prime}=t}^{T}\gamma^{t^{\prime}-t}r_{t^{\prime}}\right].$ (20)
@@
where $\delta_{t}=A_{t}^{\boldsymbol{\theta}^{\prime}}\left(\boldsymbol{s}_{t},\boldsymbol{a}_{t}\right)$. $\lambda\in(0,1]$ is a discount factor. Note that (22) aims to balance the variance and bias of the advantage estimates by leveraging information from multiple future steps while discounting their influence.
Figure 4: The framework of the MAPPO algorithm.
If there are no constraints on the policies $\pi_{\boldsymbol{\theta}}$ and $\pi_{\boldsymbol{\theta}^{\prime}}$, significant fluctuations may occur during policy updates. This is due to the limited sampling, which fails to accurately represent the probability distributions described in Lemma 1. To constrain the divergence between policies $\pi_{\boldsymbol{\theta}}$ and $\pi_{\boldsymbol{\theta}^{\prime}}$, $J^{CLIP}(\boldsymbol{\theta})$ is employed. It is defined as
@@
$=\mathbb{E}_{t}\left[\min\left(r_{t}(\boldsymbol{\theta})A_{t}^{\boldsymbol{\theta}^{\prime}}\left(\boldsymbol{s}_{t},\boldsymbol{a}_{t}\right),g\left(\epsilon,A_{t}^{\boldsymbol{\theta}^{\prime}}\left(\boldsymbol{s}_{t},\boldsymbol{a}_{t}\right)\right)\right)\right],$
@@ IV Simulation Results
In this section, the performance of the semantic-aware multi-task offloading model will be demonstrated. We provide specific implementation details, investigate the QoE performance of the proposed approach, and examine the impact of user preferences on the system.
@@ IV-A Implementation Details
Three tasks are considered: text classification, image classification, and VQA, corresponding to text, image, and multi-modal tasks, respectively. For text classification, we use the SST-2 dataset, for image classification, the CIFAR-10 dataset, and for VQA, the VQAv2 dataset is adopted. The text semantic extraction network utilizes an advanced BERT-based [45] text embedding model, while the image semantic extraction network is based on the Vision Transformer (ViT) [46]. For VQA, the encoder architecture employs BERT for linguistic processing and ViT for visual feature extraction, with cross-modal fusion and answer generation performed at the receiver. Training is conducted using the AdamW optimizer, with a learning rate of $3\times 10^{-5}$, a batch size of 50, and a weight decay of $1\times 10^{-4}$. All experiments were conducted on a platform equipped with an NVIDIA RTX 4090 and a CPU Intel i9-13900K @5.8 GHz. When the training of all semantic transmission models is complete, we can calculate the GPU consumption for basic semantic extraction $l_{n}^{SE}$ and the GPU consumption $l^{S}$ for execution at the ES.
We consider scenarios with four UEs and orthogonal sub-carriers, where each UE has a queue length of 20, containing tasks from different task types. Additional details regarding the wireless environment, GPU computational frequency, and the proposed semantic-aware MAPPO algorithm training configurations are summarized in Table II. To verify the effectiveness of the semantic-aware multi-task offloading framework, we consider the following benchmarks:
@@
Number of UEs $N$ 4
Sub-channel bandwidth $B$ 10 MHz
Task queue length $Q_{n}$ 20
@@
Transmit power $p_{n}$ range (10, 90) mW
Semantic extraction factor $\mu_{n}$ (0.1, 1)
The number of user CUDA cores 1280
Local GPU frequency $f_{n}$ range (1.5, 1.7) GHz
The number of ES CUDA cores 65536
Remote GPU frequency $f_{s}$ 2.2 GHz
@@
Energy consumption of ES $\alpha^{S}$ $1.2\times 10^{-26}$
Minimum accuracy requirements $\epsilon_{\min}$ 50%
@@
Batch size of training the MAPPO 2
@@
Learning rate $1\times 10^{-4}$
Advantage discount factor $\lambda$ 0.95
Reward discount factor $\gamma$ 0.99
@@
Policy entropy bonus weight $b_{2}$ 0.1
  • Semantic-unaware MAPPO: The benchmark transmits all data without semantic extraction during offloading, i.e. $\mu_{n}=1$. All other settings are identical to those used in semantic-aware MAPPO.
  • Semantic-aware Dueling Double DQN (D3QN): D3QN can only handle discrete action spaces, so we discretized the action space and used the same environment as semantic-aware MAPPO.
  • Semantic-unaware D3QN: This benchmark transmits all data without semantic extraction during offloading, i.e. $\mu_{n}=1$. All other settings are identical to those used in semantic-aware D3QN.
@@
All experimental results are averaged over 200 test runs to mitigate the effects of uneven task type distribution and dynamic changes in the channel.
@@ IV-B QoE Performance and Convergence
1) Convergence Performance: The preference parameters, $\omega_{t}$, $\omega_{e}$, and $\omega_{a}$ are all set to 1/3, indicating that the model has no specific preference among latency, energy consumption, and accuracy. Fig. 6 shows that the proposed MAPPO algorithm converges after approximately 100 episodes and maintains optimal performance. For the D3QN algorithm, the replay buffer size is set to 30000. Fig. 7 shows that the D3QN algorithm finally converges after 2000 episodes. The proposed MAPPO algorithm directly optimizes cumulative returns through policy gradient methods, while employing a trust region mechanism to constrain policy update magnitudes, thereby preventing training instability caused by abrupt policy changes. In contrast, the D3QN algorithm relies on TD learning and demonstrates lower learning efficiency in environments with sparse rewards or delayed feedback. Furthermore, to enhance the sampling efficiency of on-policy algorithms, the proposed MAPPO algorithm employs multi-threaded asynchronous sampling and completes state collection and synchronization during network updates. Consequently, the proposed MAPPO algorithm achieves faster convergence with greater stability compared to D3QN.
@@
Figure 7: Convergence of the D3QN reward.
2) QoE Performance: Fig. (8) reveals the relationship between QoE and the sub-carrier bandwidth for UEs. All schemes are initially trained at a sub-carrier bandwidth of 10 MHz and then tested under various bandwidth conditions. It is observed that the proposed semantic-aware MAPPO algorithm achieves the best performance across all bandwidth conditions. Comparative evaluations demonstrate 12.68% QoE improvements versus semantic-unaware MAPPO and 14.48% superiority versus semantic-unaware D3QN. Semantic-unaware methods fix $\mu_{n}=1$ to transmit all data without compression, increasing transmission and execution latency, consequently degrading QoE. The D3QN algorithm requires discrete action space approximation, which may introduce quantization errors or reduce sampling efficiency. In contrast, the proposed MAPPO algorithm directly outputs continuous action probability distributions, eliminating the need for discretization and thereby avoiding the curse of dimensionality in high-dimensional action spaces. Consequently, our method achieves superior QoE. Local execution is solely related to the local computation frequency. When calculating QoE, the local computation frequency is fixed, meaning we use the local execution’s latency, energy consumption, and accuracy to normalize the latency, energy, and accuracy obtained from other methods. Thus, the QoE of local execution is fixed at 0.5.
Figure 8: QoE with different sub-carrier bandwidths.
Figure 9: QoE with different noise powers.
Figure 10: QoE with different numbers of users.
Figure 11: QoE with different semantic extraction factor $\mu_{n}$.
Fig. (9) describes how QoE varies with channel noise intensity. All schemes are first trained under the noise intensity of 10 mW and then tested at different noise intensities. Similarly, the proposed semantic-aware MAPPO algorithm performs better than all other benchmarks. The semantic-aware algorithm dynamically selects optimal semantic extraction factors based on real-time noise intensity and channel gain, maintaining stable QoE without significant degradation even under increasing noise levels. In contrast, semantic-unaware schemes lack adaptive semantic symbol quantity regulation, resulting in substantial performance deterioration. Fig. (10) illustrates the QoE trend versus the number of users. As the number of users increases, the server must handle more computation offloading requests, leading to elevated processing latency and energy consumption, which collectively degrade overall QoE. Semantic-unaware approaches cannot leverage semantic extraction to reduce uplink latency during user scaling, ultimately forcing all users to execute tasks locally. Consequently, its QoE converges to the local computing baseline. Fig. (11) illustrates the QoE variation versus $\mu_{n}$. While trained at $\mu_{n}=0.1$, the system is evaluated across varying $\mu_{n}$. Results demonstrate that when users have fewer selectable semantic extraction factors, the inability to utilize smaller compression factors for data size reduction significantly impairs channel adaptation under poor conditions, ultimately degrading QoE.
TABLE III: Performance under different preferences
Delay (ms) Energy Consumption (J) Accuracy (%)
No Preference 0.523 0.0186 51.37
Delay Preference 0.369 0.0192 43.08
Energy Preference 3.50 0.0163 56.85
Accuracy Preference 1.67 0.0682 71.99
To further investigate the contributions of latency, energy consumption, and task performance to QoE, Fig. (12) illustrates the execution latency, energy consumption, and task accuracy. All metrics are measured under the following conditions: 10 MHz bandwidth, 2 mW noise power, and 4 users. Local task execution demonstrates longer latency and higher energy consumption, yet achieves optimal task performance. This occurs because local processing avoids performance degradation induced by semantic compression. However, the prolonged execution time incurs significant energy overhead. Semantic-unaware approaches exhibit inferior latency performance because of the transmission of raw task data without semantic extraction, resulting in extended transmission delays. However, server-side processing of offloaded tasks requires minimal time, yielding lower energy consumption compared to local execution. Owing to wireless channel effects during task offloading, semantic-unaware methods show slightly reduced task accuracy relative to local processing. The semantic-aware approach dynamically selects optimal semantic extraction factors based on real-time channel conditions, computational load, and transmission requirements, achieving the best trade-off among latency, energy efficiency, and task accuracy. Consequently, it delivers the highest QoE.
Fig. (13) presents the latency, energy consumption, and task performance across different task queue configurations. The system was initially trained on hybrid task scenarios and subsequently evaluated under four distinct test conditions: hybrid tasks, text classification, image classification, and VQA. The results demonstrate our algorithm’s strong adaptability to diverse task types. Specifically, text classification tasks exhibit the lowest latency and energy consumption while maintaining high accuracy, attributable to their smaller data volume and computational complexity. In contrast, image classification and VQA tasks require substantially more resources, resulting in higher energy expenditure and longer processing times. The hybrid task shows performance metrics that approximately equal the arithmetic mean of individual task types, confirming the algorithm’s robustness. The significant performance variations observed across task categories underscore the necessity of employing locally normalized QoE metrics for fair evaluation.
Figure 12: Performance of different metrics.
Figure 13: Performance with different task types.
IV-C QoE Preference
We can adjust the model to achieve desired results by modifying QoE preferences for different metrics. Table III shows the performance in execution latency, energy consumption, and task accuracy under varying preference settings. Specifically, we adjust the relative magnitude of $\omega_{t}$, $\omega_{e}$, and $\omega_{a}$ to set preferences for different metrics and retrain to obtain the corresponding model for each. 1) For the delay-preference model, we set $\omega_{t}=1$, $\omega_{e}=0$, and $\omega_{a}=0$. The delay-preference model tends to choose a smaller $\mu_{n}$ to reduce the amount of data transmitted. This results in a partial loss of semantic information, causing a decline in task accuracy. 2) For the energy-preference model, we set $\omega_{t}=0$, $\omega_{e}=1$ and $\omega_{a}=0$. The energy-preference model prioritizes reduction of local processing energy by favoring larger semantic extraction factors, achieving significant energy savings at the cost of substantially increased transmission latency. 3) For the accuracy-preference model, we set $\omega_{t}=0$, $\omega_{e}=0$, and $\omega_{a}=1$. The accuracy-preference model aims to maximize the accuracy of all tasks, prompting it to execute locally. Nevertheless, Local task execution exhibits significantly higher latency and energy consumption.
Fig. 14 shows the performance of latency and task accuracy under different weights by adjusting the weights assigned to latency and accuracy. As the weight for latency increases, tasks are offloaded to the ES with more powerful processing capabilities, and there tends to be a preference for a smaller semantic extraction factor, $\mu_{n}$, to maximize the compression rate. This results in a decrease in the accuracy of tasks executed by the ES. Fig. 15 presents the performance of energy consumption and task accuracy under different weights. A significant jump occurs at (0.8, 0.2) when the weight for accuracy increases, as tasks are executed locally, causing a substantial increase in energy consumption and a corresponding rise in accuracy. Fig. 16 describes the performance of energy consumption and latency under varying weights. When the weight for latency increases beyond (0.3, 0.7), latency begins to decrease, and energy consumption increases. This is due to an increase in the number of tasks offloaded to the ES, and users tend to choose a smaller semantic extraction factor, $\mu_{n}$, to reduce data transmission volume. However, a smaller $\mu_{n}$ means that users need more computational resources to extract more compact semantic information.